OpenStudy (anonymous):

Here is limit question\[\lim_{x \rightarrow 0}\left( \frac{ (9^{x}-1)(4^{x}-1) }{ \sqrt{2}-\sqrt{1+\cos(x)}} \right)\]

OpenStudy (anonymous):

i hate this type of limit questions

hartnn (hartnn):

divide numerator and denom. by x^2

hartnn (hartnn):

lim (a^x-1)/x=ln a

hartnn (hartnn):

multiply and divide by conjugate : \(\sqrt2+\sqrt{1+cos x}\)

OpenStudy (anonymous):

not that way @hartnn

hartnn (hartnn):

in denominator, 1-cos x/x^2 remains

hartnn (hartnn):

which is half, and u had ln9.ln4. final answer ln9.ln/root 2 ??

hartnn (hartnn):

ln9.ln4/root 2 ??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Let me give you the answer first\[8sqrt{2}(ln 3)(ln 2) \]

hartnn (hartnn):

L'Hopital's ?

OpenStudy (anonymous):

nope

hartnn (hartnn):

ln9.ln4/root 2 is exactly same as 8sqrt2(ln3)(ln2)

hartnn (hartnn):

i was correct, just needed to simplify

OpenStudy (anonymous):

you are not @ hartnn

hartnn (hartnn):

ln9=2ln 3 ln4 =2ln2 ln9.ln4/root 2 = 4ln 2.ln3 / root 2 = 8 ln2.ln3 *root2 got it ?

hartnn (hartnn):

nopes, my expression is exactly simplified to your answer.

OpenStudy (anonymous):

you are incorrect while you rationalize sqrt 2

hartnn (hartnn):

can't find how 8 comes......yes, i m getting 2root2 instead of 8root2

OpenStudy (anonymous):

then...

hartnn (hartnn):

let me write it properly again.

hartnn (hartnn):

\(\huge\frac{4ln3.ln2*2\sqrt2}{1/2}\) i made the mistake of writing 2root 2 in denom. actually its in numerator. now i m getting 16

OpenStudy (nipunmalhotra93):

I'm getting 16 instead of 8 lol

OpenStudy (nipunmalhotra93):

@hartnn yea..

OpenStudy (anonymous):

Right, Here is how you solve it @hartnn

hartnn (hartnn):

me too getting 16

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\left( \frac{ (9^{x}-1)(4^{x}-1) }{ \sqrt{2}-\sqrt{1+\cos(x)}} \right)\] \[\lim_{x \rightarrow 0}\left( \frac{ (9^{x}-1)(4^{x}-1)times(sqrt{2}+sqrt{1+cos(x)}) }{ 2-1-cos(x)} \right) \] \[\lim_{x \rightarrow 0}\left( \frac{ (9^{x}-1)(4^{x}-1)(sqrt{2}+sqrt{1+cos(x)}) }{ 1-sqrt{1-sin^{2}x }} \right) \] \[\lim_{x \rightarrow 0}( \frac{ (9^{x}-1)(4^{x}-1)(sqrt{2}+sqrt{1+cos(x)})times(1+sqrt{1-sin^{2}x)} }{ sin^{2}x } ) \] \[\lim_{x \rightarrow 0}( (\frac{ x }{ sin(x) })^{2}times(\frac{ 9^{x}-1 }{ x })times(\frac{ 4^{x}-1 }{ x })times(sqrt{2}+sqrt{1+cos(x)})times(1+sqrt{1-sin^{2}x}))\] \[ (1)^{2}times(ln 9)times(ln 4)times(2sqrt{2})times(1) \] \[8sqrt{2}(ln 3)(ln 2) \]

OpenStudy (anonymous):

I am closing this

hartnn (hartnn):

wait how u getting 8, it should be 16

OpenStudy (anonymous):

See the steps

hartnn (hartnn):

u forgot one of the '2's for 1+sqrt(1-sin^2 x)

OpenStudy (anonymous):

Ya,ya ...I forget it thanks it 16.

hartnn (hartnn):

u wrote it as times 1

hartnn (hartnn):

it should be times 2

OpenStudy (anonymous):

It is true